# thanks: The first two authors contributed equally

### v1

@@ Abstract
The Inspiral Merger Ringdown Consistency Test (IMRCT) is one among a battery of tests of general relativity (GR) employed by the LIGO-Virgo-KAGRA (LVK) collaboration. It is used to search for deviations from GR in detected gravitational waves (GWs) from compact binary coalescences (CBCs) in a model-agnostic way. The test compares source parameter estimates extracted independently from the inspiral and post-inspiral portions of the CBC signals and, therefore, crucially relies on the accurate modeling of the waveform. Current implementations of the IMRCT routinely use quasicircular waveforms, under the assumption that the residual eccentricity of the binary when the emitted GWs enter the frequency band of the LVK detector network will be negligible. In this work, we perform a detailed study to investigate the typical magnitudes of this residual eccentricity that could potentially lead to spurious violations of the IMRCT. To that end, we conduct injection campaigns for a range of eccentricities and recover with both quasicircular and eccentric waveforms. We find that an eccentric GW signal from a GW150914-like system with $e_{\text{gw}}\gtrsim$ 0.04 at an orbit averaged frequency $\langle f_{\text{ref}}\rangle=$ 25 Hz breaks the IMRCT if recovered with quasicircular waveforms at $\gtrsim 68\%$ confidence. The violation becomes more severe ($\gtrsim 90\%$ confidence) for $e_{\text{gw}}=$ 0.055 at $\langle f_{\text{ref}}\rangle=$ 25 Hz. On the other hand, when eccentric waveforms are used, the IMRCT remains intact for all eccentricities considered. We also briefly investigate the effect of the magnitude and orientation (aligned/antialigned) of the component spins of the binary on the extent of the spurious violations of the IMRCT. Our work, therefore, demonstrates the need for accurate eccentric waveform models in the context of tests of GR.
@@ I Introduction
The detected GW signals, especially those pertaining to BBH mergers, come from the last few orbits of the inspiral, as well as the merger and ringdown phases (see, e.g., Sathyaprakash and Schutz (2009)). These GWs are well-suited to probe general relativity (GR) in the strong-field regime, unlike most tests of GR that use electromagnetic waves, which typically probe the weak field regime [^1] A notable exception to this are the results of the Event Horizon Telescope (EHT) Akiyama et al. (2019), which imaged the shadow of the supermassive black hole at the center of the M31 galaxy and enabled EM-based strong-field tests of GR Psaltis (2019)..
The LIGO-Virgo-KAGRA (LVK) collaboration  Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2021) has conducted a suite of tests of GR across O1, O2, and O3 Abbott et al. (2021b). This includes222but is not limited to a model-agnostic residuals test, which subtracts out the best-matched GR-modeled CBC waveform from the data containing the GW signal, and checks if the statistical properties of the residual are consistent with noise Abbott et al. (2016, 2019a, 2020b); a test that probes the inspiral evolution of the CBC by searching for deviations in Post-Newtonian (PN) parameters of the GW signal Blanchet and Sathyaprakash (1995, 1994); Arun et al. (2006a, b); Yunes and Pretorius (2009); Mishra et al. (2010); Li et al. (2012a, b); an inspiral-merger-ringdown consistency test (IMRCT) that compares and assesses the consistency between the low and high-frequency portions of the CBC signals Ghosh et al. (2016, 2018); and propagation tests of GWs that compare their speed with respect to the speed of light Abbott et al. (2019b) as well as any modulations in the waveform due to finite-graviton-mass-driven velocity dispersion Will (1998).
In this work, we restrict our attention to the IMRCT. The IMRCT compares the two-dimensional joint posterior on the final mass and spin of the merged binary estimated independently from the inspiral (low-frequency) and post-inspiral (high-frequency) portions of the waveform Ghosh et al. (2016, 2018). A sufficiently large deviation from GR should result in the difference distribution evaluated from these 2D posteriors to deviate from the origin (zero). However, it is conceivable that waveforms that do not accurately capture the physics of the CBC could lead to biased posteriors. We propose to study the resulting systematics in the IMRCT. In particular, we investigate how the neglect of eccentricity in the waveform models could cause spurious violations of this test.
@@
On the other hand, certain formation channels, such as those pertaining to dense stellar environments, could produce binaries whose eccentricities are non-negligible when they enter the LVK frequency band Mapelli (2020). Dynamical encounters in such environments could harden the binary so that GW emission cannot completely carry away the eccentricity by the time the GWs enter the LVK’s frequency band. Moreover, in AGN disks, eccentricity can be amplified by binary-single encounters Samsing et al. (2022); Tagawa et al. (2021), while the presence of a third object in the vicinity of the binary could similarly boost the eccentricity via the Kozai-Lidov mechanism Kozai (1962); Lidov (1962); Naoz (2016); Antonini et al. (2017); Randall and Xianyu (2018); Yu et al. (2020); Bartos et al. (2023). Such eccentricity-enhancing processes could also leave a residual in-band eccentricity detectable by the LVK network. Neglecting eccentricity of such signals in (PN) parametrized tests has already been shown to lead to spurious violations of GR both at the individual event level Saini et al. (2022) and population level inference Saini et al. (2023). We here focus only on the effect of neglect of eccentricity on the IMRCT.
@@
To assess the impact of neglecting eccentricity in waveform models on the IMRCT, we consider a set of synthetic eccentric GW signals. We inject these in the most probable realization of (zero-mean) Gaussian noise, the so-called “zero-noise” realization, assuming an O4-like noise power-spectral density (PSD). We then perform the IMRCT on each of these signals, using quasicircular templates. We find that the IMRCT is indeed violated with quasicircular templates, as we increase the injected eccentricity, for a set of masses and spins where the IMRCT is known not to suffer from other systematics. This violation progressively increases, from moderate to severe, with increasing eccentricity. To ascertain that no intrinsic waveform systematics are contributing to this bias, we repeat the IMRCT using eccentric waveform model for recovery. We indeed find that the IMRCT is no longer violated.
@@
Past work, using an approximate Fisher Matrix Analysis (FMA) Cutler and Flanagan (1994) and the Cutler-Valisneri framework Cutler and Vallisneri (2007), pointed out that the IMRCT could be violated due to neglect of eccentricity Bhat et al. (2023). Another work also found spurious violations of the IMRCT using available numerical-relativity waveforms as injections, while performing the IMRCT with quasicircular phenomenological waveform models Narayan et al. (2023). Our work complements and expands on both of these studies in the following way. We perform Bayesian parameter inference runs using large-scale nested sampler and, therefore, have more realistic estimate of the bias compared to those in provided in Bhat et al. (2023). Moreover, unlike in Bhat et al. (2023), we do not neglect the eccentricity in the post-inspiral part of the signal which may lead to over-estimation or under-estimation of the bias. Also, because we employ an eccentric waveform model TEOBResumS-Dalí Nagar et al. (2018, 2021), we have two advantages over the results in  Narayan et al. (2023). We use the same waveform model to create the injection and then to recover it. Therefore, we rule out any bias originating due to the waveform systematics which may be present if the recovery template is not as accurate as the injected NR waveforms. Also, we have the freedom to create injections at arbitrary values of eccentricities allowing us to do an injection study on a finer grid of eccentricity values. This, in turn, allows us to identify more accurately at what eccentricity the IMRCT bias becomes severe. Similarly, we examine the sensitivity of the bias for a given eccentricity on other parameters like the magnitude of the component spins and their orientation (aligned/antialigned) with respect to the orbital angular momentum of the binary. Moreover, we ascertain that the IMRCT is indeed recovered as expected, when eccentricity is included, for all eccentricities considered. We make contact with both these works to compare the eccentricity at which the IMRCT is violated, using a generalised definition of eccentricity that is independent of waveform models Shaikh et al. (2023).
@@ II.1 Gravitational-wave parameter estimation
In GR, GWs have two polarizations, $h_{+}$ and $h_{\times}$. The phase evolution of these polarizations depends on the intrinsic parameters of the binary, viz. the masses $m_{1},m_{2}$ of the individual components  [^3] Strictly, the masses that determine the shape of the GW signal measured at the detector are the so-called “detector-frame” masses, which are simply the masses multiplied by $(1+z)$, where $z$ is the cosmological redshift., and the spin vectors of these components $\mathbf{S}_{1},\mathbf{S}_{2}$. The amplitude of the polarizations depend on the inclination of the binary with respect to the line of sight, $\iota$, luminosity distance $d_{L}$, the phase $\phi_{c}$ and time $t_{c}$ at coalescence. In addition to the above mentioned intrinsic parameters, we need two more parameters – eccentricity $e$ and mean anomaly $l$  [^4] The angle variable that describes the position of the compact objects on the orbit can be also described by other equivalent variables like eccentric anomaly or true anomaly. Clarke et al. (2022) – at a given reference frequency to accurately describe GWs from a binary in eccentric orbit. Also, the GW waveform shape is more sensitive to the ‘chirpmass’ $\mathcal{M}_{c}=m_{1}^{3/5}m_{2}^{3/5}/(m_{1}+m_{2})^{1/5}$ compared to individual masses and often it is more practical to use $\mathcal{M}_{c}$ and the mass ratio $q=m_{1}/m_{2}$ instead of $m_{1}$ and $m_{2}$ for parameter estimation.
@@ II.2 The inspiral-merger-ringdown consistency test
In this work, we restrict ourselves to non-precessing waveforms, i.e, waveforms produced by binaries whose spin-vectors are aligned or antialigned with the orbital angular momentum vector $\mathbf{L}$ ($z$-axis). Moreover, we only consider binaries with mass-ratios $q$ that are sufficiently small ($q=m_{1}/m_{2}\sim 1$). We therefore only need to use the leading order $(2,\pm 2)$ multipole, since higher multipoles for the types of systems we consider will be suppressed.
The choice of the transition frequency that separates the inspiral and post-inspiral portions of the waveform is not unique. However, all implementations of the IMRCT use some approximation of the frequency corresponding to the innermost stable circular orbit (ISCO) of the binary. The LVK estimates the ISCO using the medians of intrinsic parameters $\theta$ extracted from the full IMR waveform. Narayan et al Narayan et al. (2023) use the medians of the final mass and spin of the merged binary estimated from the full IMR waveform, similar to what was originally done in Ghosh et al. (2016, 2018). However, as pointed out in Ghosh et al. (2016, 2018); Narayan et al. (2023), the IMRCT is largely insensitive to moderate variations in the choice of the transition frequency. In this work, we use the exact values of $\vec{\theta}$ (i.e, the injected values), as well as the numerical relativity fits Hofmann et al. (2016); Healy and Lousto (2017); Jiménez-Forteza et al. (2017), to determine a point-estimate of the final mass and spin, from which the ISCO frequency is evaluated.
@@ II.3 A standardized definition of eccentricity
In GR, there is no universally agreed-upon definition of eccentricity. Consequently, different waveform models adopt custom internal definitions of eccentricity, leading to model-dependent inferences about eccentricity. To mitigate ambiguity in the inferred value of eccentricity from gravitational wave data analysis, there is a growing effort to standardize the definition of eccentricity. This standardization involves measuring eccentricity from a gauge-independent quantity, such as the gravitational waveform modes Ramos-Buades et al. (2020); Islam et al. (2021); Bonino et al. (2023); Shaikh et al. (2023).
@@
After computing $e_{\omega_{22}}$ from $\mathpzc{h}_{22}$, $e_{\text{gw}}$ is determined through the following transformation to ensure that $e_{\text{gw}}$ correctly exhibits the Newtonian limit Ramos-Buades et al. (2022a).
@@
Since eccentricity decays via gravitational wave radiation, we must specify a reference frequency $\langle f_{\text{ref}}\rangle$ where the value of $e_{\text{gw}}$ is measured. Moreover, eccentricity also induces modulations in the instantaneous frequency $\omega_{22}$, we use $\langle\omega_{22}\rangle$ (by taking the orbital average of the instantaneous $\omega_{22}$, see Section II.E.2 of Shaikh et al. (2023) for more details) instead of $\omega_{22}$ for defining the reference frequency $\langle f_{\text{ref}}\rangle$. Presenting our results with a standardized definition facilitates one-to-one comparisons with similar studies in the past and enables future studies to do the same with the results in this work.

### v2

@@ Abstract
The Inspiral Merger Ringdown Consistency Test (IMRCT) is one among a battery of tests of general relativity (GR) employed by the LIGO-Virgo-KAGRA (LVK) collaboration. It is used to search for deviations from GR in detected gravitational waves (GWs) from compact binary coalescences (CBCs) in a model-agnostic way. The test compares source parameter estimates extracted independently from the inspiral and post-inspiral portions of the CBC signals and, therefore, crucially relies on the accurate modeling of the waveform. Current implementations of the IMRCT routinely use quasicircular waveforms, under the assumption that the residual eccentricity of the binary when the emitted GWs enter the frequency band of the LVK detector network will be negligible. In this work, we perform a detailed study to investigate the typical magnitudes of this residual eccentricity that could potentially lead to spurious violations of the IMRCT. To that end, we conduct injection campaigns for a range of eccentricities and recover with both quasicircular and eccentric waveforms. We find that an eccentric GW signal from a GW150914-like system with eccentricity $e_{\text{gw}}\gtrsim$ 0.04 at an orbit averaged frequency $\langle f_{\text{ref}}\rangle=$ 25 Hz breaks the IMRCT if recovered with quasicircular waveforms at $\gtrsim 68\%$ confidence. The violation becomes more severe ($\gtrsim 90\%$ confidence) for $e_{\text{gw}}=$ 0.055 at $\langle f_{\text{ref}}\rangle=$ 25 Hz. On the other hand, when eccentric waveforms are used, the IMRCT remains intact for all eccentricities considered. We also briefly investigate the effect of the magnitude and orientation (aligned/antialigned) of the component spins of the binary on the extent of the spurious violations of the IMRCT. Our work, therefore, demonstrates the need for accurate eccentric waveform models in the context of tests of GR.
@@ I Introduction
The detected GW signals, especially those pertaining to BBH mergers, come from the last few orbits of the inspiral, as well as the merger and ringdown phases (see, e.g., Sathyaprakash and Schutz (2009)). These GWs are well-suited to probe general relativity (GR) in the strong-field regime, unlike most tests of GR that use electromagnetic waves, which typically probe the weak field regime [^1] A notable exception to this are the results of the Event Horizon Telescope (EHT) Akiyama et al. (2019), which imaged the shadow of the supermassive black hole at the center of the M31 galaxy and enabled electromagnetic-based strong-field tests of GR Psaltis (2019)..
The LIGO-Virgo-KAGRA (LVK) collaboration  Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2021) has conducted a suite of tests of GR across O1, O2, and O3 Abbott et al. (2021b). This includes222but is not limited to
  • A model-agnostic residuals test, which subtracts out the best-matched GR-modeled CBC waveform from the data containing the GW signal, and checks if the statistical properties of the residual are consistent with noise Abbott et al. (2016, 2019a, 2020b)
  • A test that probes the inspiral evolution of the CBC by searching for deviations in Post-Newtonian (PN) parameters of the GW signal Blanchet and Sathyaprakash (1995, 1994); Arun et al. (2006a, b); Yunes and Pretorius (2009); Mishra et al. (2010); Li et al. (2012a, b)
  • An inspiral-merger-ringdown consistency test (IMRCT) that compares and assesses the consistency between the low and high-frequency portions of the CBC signals Ghosh et al. (2016, 2018)
  • Propagation tests of GWs that compare their speed with respect to the speed of light Abbott et al. (2019b) as well as any modulations in the waveform due to finite-graviton-mass-driven velocity dispersion Will (1998).
In this work, we restrict our attention to the IMRCT. The IMRCT compares the two-dimensional joint posterior on the final mass and spin of the merged binary estimated independently from the inspiral (low-frequency) and post-inspiral (high-frequency) portions of the waveform Ghosh et al. (2016, 2018). A sufficiently large deviation from GR should result in the difference distribution between final mass and final spin evaluated from these 2D posteriors to deviate from the origin (zero). However, it is conceivable that waveforms that do not accurately capture the physics of the CBC could lead to biased posteriors. We propose to study the resulting systematics in the IMRCT. In particular, we investigate how the neglect of eccentricity in the waveform models could cause spurious violations of this test.
@@
On the other hand, certain formation channels, such as those pertaining to dense stellar environments, could produce binaries whose eccentricities are non-negligible when they enter the LVK frequency band Mapelli (2020). Dynamical encounters in such environments could harden the binary so that GW emission cannot completely carry away the eccentricity by the time the GWs enter the LVK’s frequency band Samsing (2018); Zevin et al. (2019); Romero-Shaw et al. (2022). Moreover, in active-galactic-nuclei disks, eccentricity can be amplified by binary-single encounters Samsing et al. (2022); Tagawa et al. (2021), while the presence of a third object in the vicinity of the binary could similarly boost the eccentricity via the Kozai-Lidov mechanism Kozai (1962); Lidov (1962); Naoz (2016); Antonini et al. (2017); Randall and Xianyu (2018); Yu et al. (2020); Bartos et al. (2023). Such eccentricity-enhancing processes could also leave a residual in-band eccentricity detectable by the LVK network. Neglecting eccentricity of such signals in (PN) parametrized tests has already been shown to lead to spurious violations of GR both at the individual event level Saini et al. (2022) and population level inference Saini et al. (2023). We here focus only on the effect of neglect of eccentricity on the IMRCT.
@@
To assess the impact of neglecting eccentricity in waveform models on the IMRCT, we consider a set of synthetic eccentric GW signals. We inject these in the most probable realization of (zero-mean) Gaussian noise, the so-called “zero-noise” realization [^3] Here zero-noise realization refers to the most-probable realization of Gaussian noise with zero mean and the parameter uncertainties obtained using this noise-realization provide lower bound to the estimates obtained using the non-zero real noise-realization., assuming an O4-like noise power-spectral density (PSD) Abbott et al. (2018). We then perform the IMRCT on each of these signals, using quasicircular templates. We find that the IMRCT is indeed violated with quasicircular templates, as we increase the injected eccentricity, for a set of masses and spins where the IMRCT is known not to suffer from other systematics. This violation progressively increases, from moderate to severe, with increasing eccentricity. To ascertain that no intrinsic waveform systematics are contributing to this bias, we repeat the IMRCT using eccentric waveform model for recovery. We indeed find that the IMRCT is no longer violated.
@@
Past work, using an approximate Fisher Matrix Analysis (FMA) Cutler and Flanagan (1994) and the Cutler-Vallisneri framework Cutler and Vallisneri (2007), pointed out that the IMRCT could be violated due to neglect of eccentricity Bhat et al. (2023). Another work also found spurious violations of the IMRCT using available numerical-relativity waveforms as injections, while performing the IMRCT with quasicircular phenomenological waveform models Narayan et al. (2023). Our work complements and expands on both of these studies in the following way. We perform Bayesian parameter inference runs using large-scale nested sampler and, therefore, have more realistic estimate of the bias compared to those in provided in Bhat et al. (2023). Moreover, unlike in Bhat et al. (2023), we do not neglect the eccentricity in the post-inspiral part of the signal which may lead to over-estimation or under-estimation of the bias. Also, because we employ an eccentric waveform model TEOBResumS-Dalí Nagar et al. (2018, 2021a), we have two advantages over the results in  Narayan et al. (2023). We use the same waveform model to create the injection and then to recover it. Therefore, we rule out any bias originating due to the waveform systematics which may be present if the recovery template is not as accurate as the injected NR waveforms. Also, we have the freedom to create injections at arbitrary values of eccentricities allowing us to do an injection study on a finer grid of eccentricity values. This, in turn, allows us to identify more accurately at what eccentricity the IMRCT bias becomes severe. Similarly, we examine the sensitivity of the bias for a given eccentricity on other parameters like the magnitude of the component spins and their orientation (aligned/antialigned) with respect to the orbital angular momentum of the binary. Moreover, we ascertain that the IMRCT is indeed recovered as expected, when eccentricity is included, for all eccentricities considered. We make contact with both these works to compare the eccentricity at which the IMRCT is violated, using a generalised definition of eccentricity that is independent of waveform models Shaikh et al. (2023).
@@ II.1 Gravitational-wave parameter estimation
In GR, GWs have two polarizations, $h_{+}$ and $h_{\times}$. The phase evolution of these polarizations depends on the intrinsic parameters of the binary, viz. the masses $m_{1},m_{2}$ of the individual components  [^4] Strictly, the masses that determine the shape of the GW signal measured at the detector are the so-called “detector-frame” masses, which are simply the masses multiplied by $(1+z)$, where $z$ is the cosmological redshift., and the spin vectors of these components $\mathbf{S}_{1},\mathbf{S}_{2}$. The amplitude of the polarizations depend on the inclination of the binary with respect to the line of sight, $\iota$, luminosity distance $d_{L}$, the phase $\phi_{c}$ and time $t_{c}$ at coalescence. In addition to the above mentioned intrinsic parameters, we need two more parameters – eccentricity $e$ and mean anomaly $l$  [^5] The angle variable that describes the position of the compact objects on the orbit can be also described by other equivalent variables like eccentric anomaly or true anomaly. Clarke et al. (2022) – at a given reference frequency to accurately describe GWs from a binary in eccentric orbit. Also, the GW waveform shape is more sensitive to the ‘chirpmass’ $\mathcal{M}_{c}=m_{1}^{3/5}m_{2}^{3/5}/(m_{1}+m_{2})^{1/5}$ compared to individual masses and often it is more practical to use $\mathcal{M}_{c}$ and the mass ratio $q=m_{1}/m_{2}$ instead of $m_{1}$ and $m_{2}$ for parameter estimation (PE) Cutler and Flanagan (1994).
@@ II.2 The inspiral-merger-ringdown consistency test
In this work, we restrict ourselves to non-precessing waveforms, i.e, waveforms produced by binaries whose spin-vectors are aligned or antialigned with the orbital angular momentum vector $\mathbf{L}$ ($z$-axis). Moreover, we only consider binaries with mass-ratios $q$ that are close to 1 ($q=m_{1}/m_{2}\sim 1$). We therefore only need to use the leading order $(2,\pm 2)$ multipole, since higher multipoles for the types of systems we consider will be suppressed.
The choice of the transition frequency that separates the inspiral and post-inspiral portions of the waveform is not unique. However, all implementations of the IMRCT use some approximation of the frequency corresponding to the innermost stable circular orbit (ISCO) of the binary. The LVK estimates the ISCO using the medians of intrinsic parameters $\theta$ extracted from the full inspiral-merger-ringdown (IMR) waveform. Narayan et al Narayan et al. (2023) use the medians of the final mass and spin of the merged binary estimated from the full IMR waveform, similar to what was originally done in Ghosh et al. (2016, 2018). However, as pointed out in Ghosh et al. (2016, 2018); Narayan et al. (2023), the IMRCT is largely insensitive to moderate variations in the choice of the transition frequency. In this work, we use the exact values of $\vec{\theta}$ (i.e, the injected values), as well as the numerical relativity fits Hofmann et al. (2016); Healy and Lousto (2017); Jiménez-Forteza et al. (2017), to determine a point-estimate of the final mass and spin, from which the ISCO frequency is evaluated.
@@ II.3 A standardized definition of eccentricity
In GR, there is no universally agreed-upon definition of eccentricity. Consequently, different waveform models adopt custom internal definitions of eccentricity, leading to model-dependent inferences about eccentricity. To mitigate ambiguity in the inferred value of eccentricity from gravitational wave data analysis, there is a growing effort to standardize the definition of eccentricity. This standardization involves measuring eccentricity from a gauge-independent quantity, such as the gravitational waveform modes Mora and Will (2002); Ramos-Buades et al. (2020); Islam et al. (2021); Nagar et al. (2021a); Ramos-Buades et al. (2022a); Bonino et al. (2023); Ramos-Buades et al. (2022b); Shaikh et al. (2023).
@@
After computing $e_{\omega_{22}}$ from $\mathpzc{h}_{22}$, $e_{\text{gw}}$ is determined through the following transformation to ensure that $e_{\text{gw}}$ correctly exhibits the Newtonian limit Ramos-Buades et al. (2022b).
@@
Since eccentricity decays via gravitational wave radiation, we must specify a reference frequency where the value of $e_{\text{gw}}$ is measured. Because eccentricity induces modulations in the instantaneous frequency $\omega_{22}$, we obtain a monotonic orbit-averaged $\langle\omega_{22}\rangle$ by taking the orbital average of the instantaneous $\omega_{22}$ (see Section II.E.2 of Shaikh et al. (2023) for more details) Between any two consecutive extrema (local maxima (pericenters, $t^{\text{p}}$) or minima (apocenters, $t^{\text{a}}$) in the instantaneous $\omega_{22}$) $t^{\text{X}}_{i}$ and $t^{\text{X}}_{i+1}$ one can define
$\langle\omega_{22}\rangle_{i}^{\text{X}}$ $=\frac{1}{t^{\text{X}}_{i+1}-t^{\text{X}}_{i}}\,\int_{t^{\text{X}}_{i}}^{t^{\text{X}}_{i+1}}\omega_{22}(t)\,\mathrm{d}t$
$=\frac{\phi_{22}(t^{\text{X}}_{i+1})-\phi_{22}(t^{\text{X}}_{i})}{t^{\text{X}}_{i+1}-t^{\text{X}}_{i}},$ (12)
and associate $\langle\omega_{22}\rangle_{i}^{\text{X}}$ with the midpoint between $t^{\text{X}}_{i}$ and $t^{\text{X}}_{i+1}$:
$\langle t\rangle^{X}_{i}=\frac{1}{2}\left(t^{\text{X}}_{i}+t^{\text{X}}_{i+1}\right).$ (13)
Applying this procedure to all consecutive pairs of pericenter times, one can obtain the set $\big{\{}\,\left(\langle t\rangle^{\text{p}}_{i},\langle\omega_{22}\rangle_{i}^{\text{p}}\right)\,\big{\}}$ and similarly, the set $\left\{\,\big{(}\langle t\rangle^{\text{a}}_{i},\langle\omega_{22}\rangle_{i}^{\text{a}}\big{)}\,\right\}$ for apocenter times. Finally, taking the union of these two datasets, one can build an interpolant in time to obtain $\langle\omega_{22}\rangle(t)$ and use $\langle f_{\text{ref}}\rangle(t_{\text{ref}})\equiv\langle\omega_{22}\rangle(t_{\text{ref}})/(2\pi)$ for defining the reference frequency $\langle f_{\text{ref}}\rangle$ corresponding to a reference time $t_{\text{ref}}$.
Presenting our results with a standardized definition facilitates one-to-one comparisons with similar studies in the past and enables future studies to do the same with the results in this work.